OpenStudy (anonymous):

There are 15 technicians and 11 chemists working in a research laboratory. In how many ways could they form a 5-member safety committee if the committee must have exactly one technician?

OpenStudy (anonymous):

25C5

OpenStudy (anonymous):

11+15 = 26 (-1) technician = 26-1 = 25 (25)! / [(25-4)! (4)!

OpenStudy (anonymous):

12650

OpenStudy (anonymous):

that is not right.

OpenStudy (anonymous):

I realized i have not learned this part of the lesson yet, and my teacher is going to be going over it tomorrow

OpenStudy (anonymous):

btw the answer is 15C1*11C4=4950

OpenStudy (anonymous):

Oh my bad . I was thinking permutation and applied combination. Thanks for pointing out.

OpenStudy (anonymous):

No problem, we all make mistakes ;)